OpenStudy (anonymous):

solve the system of equation: 3x+6y=14, x+2y=3 a.(-7,5) b. (0,0) c.(5,-7) d.no solution

OpenStudy (anonymous):

first, you move 2y to the right side of the equation -> x = 3-2y; substitute the x into the first equation, and check whether the new equation makes sense or not. if not, then there is no solution

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

wait no you move 2y to the beginnin. and 3x right

OpenStudy (anonymous):

if you are using substitution, the order doesn't matter. if you want to use y to substitute: y = (3-x)/2, and substitute into 3x+6y=14 which will become 3x + 6((3-x)/2) = 14